# Foundation Models for Time Series Analysis: A Tutorial and Survey

### v2

@@ 5.3.2. Multi-modality
However, the single-modal methods may not encapsulate the full picture for several challenging downstream tasks in finance (Chen et al., 2023f; Yu et al., 2023) and healthcare domains (Liu et al., 2023b; Li et al., 2023). To cope with this issue, there have been initiatives towards developing multi-modal, task-oriented foundation models, where additional information provides useful information to enhance the model capability. In Chen et al., an external ChatGPT is queried to construct an evolving graph structure representing companies, based on the analysis of news headlines at specific time steps. As such, the inferred graph and stock prices are fed into the time series model (that uses GNN and LSTM for information propagation) to generate stock price movement predictions. Another example in healthcare also demonstrated the effectiveness of multi-modal medical context modeling, which aligns the embedding of ECG (Electrocardiogram) and corresponding medical text reports under a self-supervised contrastive learning framework and performs ECG classification. In general multi-modal time series analysis, similar cross-modality alignment strategies (e.g., contrastive learning (Sun et al., 2023), reprogramming (Jin et al., 2023b), token-wise prompting (Liu et al., 2024b)) are adopted, where the multi-modal inputs are often the textual description of datasets and pre-training word embedding from LLMs. As a notable example, Time-LLM (Jin et al., 2023b) introduces a reprogramming framework that aligns the language knowledge from pre-trained word embedding and time series information via linear projection and multi-head attention, where the handcrafted dataset descriptions are also used to quired text token embeddings as prompts, which further enhances the embedding space and informs the LLM to comprehend the task contexts. As such, utilizing multi-modal data facilitates the repurpose of the existing LLM into time series forecasters without additional computational costs.
6. Future Direction
@@
7. Conclusion
The rapid development of AI foundation models has revolutionized the research fields in different domains. In this survey, we provide a comprehensive and updated review of foundation models specifically designed for time series analysis. A novel taxonomy is proposed from a methodology-centric perspective by classifying FMs based on key components including model architecture, pre-training technique, adaptation technique, and data modality. Our survey facilitates understanding the underlying mechanism of applying the foundation models to time series. Furthermore, we believe that the consolidation of the latest advancements, as well as the potential future direction, can inspire more innovative works within the field of time series analysis.
Appendix A Appendix
We list the summary of representative foundation models tailored for time series data modeling in Table 2.

### v3

@@ 5.3.2. Multi-modality
However, the single-modal methods may not encapsulate the full picture for several challenging downstream tasks in finance (Chen et al., 2023f; Yu et al., 2023) and healthcare domains (Liu et al., 2023b; Li et al., 2023). To cope with this issue, there have been initiatives towards developing multi-modal, task-oriented FMs, where additional information provides useful information to enhance the model capability. In Chen et al., an external ChatGPT is queried to construct an evolving graph structure representing companies, based on the analysis of news headlines at specific time steps. As such, the inferred graph and stock prices are fed into the time series model (that uses GNN and LSTM for information propagation) to generate stock price movement predictions. Another example in healthcare also demonstrated the effectiveness of multi-modal medical context modeling, which aligns the embedding of ECG (Electrocardiogram) and corresponding medical text reports under a self-supervised contrastive learning framework and performs ECG classification. In general multi-modal time series analysis, similar cross-modality alignment strategies (e.g., contrastive learning (Sun et al., 2023), reprogramming (Jin et al., 2023b), token-wise prompting (Liu et al., 2024b)) are adopted, where the multi-modal inputs are often the textual description of datasets and pre-training word embedding from LLMs. As a notable example, Time-LLM (Jin et al., 2023b) introduces a reprogramming framework that aligns the language knowledge from pre-trained word embedding and time series information via linear projection and multi-head attention, where the handcrafted dataset descriptions are also used to quired text token embeddings as prompts, which further enhances the embedding space and informs the LLM to comprehend the task contexts. As such, utilizing multi-modal data facilitates the repurpose of the existing LLM into time series forecasters without additional computational costs.
6. Conclusion
The rapid development of FMs has revolutionized the research fields in different domains. In this survey, we provide a comprehensive and updated review of FMs specifically designed for time series analysis. A novel taxonomy is proposed from a methodology-centric perspective by classifying FMs based on key components including model architecture, pre-training technique, adaptation technique, and data modality. Our survey facilitates understanding the underlying mechanism of applying the FMs to time series. Furthermore, we believe that the consolidation of the latest advancements, as well as the potential future direction (see Appendix), can inspire more innovative works within the field of time series analysis.
Acknowledgements
This work is mainly supported by the Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0620). It is also funded by Guangzhou Municipal Science and Technology Project 2023A03J0011.
Appendix A Appendix